### v1

# MnLargeSymbols’164 MnLargeSymbols’171

Abstract
Modern deep neural networks often require distributed training with many workers due to their large size. As worker numbers increase, communication overheads become the main bottleneck in data-parallel minibatch stochastic gradient methods with per-iteration gradient synchronization. Local gradient methods like Local SGD reduce communication by only syncing after several local steps. Despite understanding their convergence in i.i.d. and heterogeneous settings and knowing the importance of batch sizes for efficiency and generalization, optimal local batch sizes are difficult to determine. We introduce adaptive batch size strategies for local gradient methods that increase batch sizes adaptively to reduce minibatch gradient variance. We provide convergence guarantees under homogeneous data conditions and support our claims with image classification experiments, demonstrating the effectiveness of our strategies in training and generalization.
1 Introduction
Recent advances in deep learning, particularly large language models (LLMs) with billions of parameters, necessitate efficient distributed training on many GPUs or TPUs. Data-parallel optimization [83] simplifies this by having each worker compute and average batch gradients from different i.i.d.  data batches, followed by a global model update. Despite its efficiency with numerous workers, this method faces challenges such as costly gradient synchronization and reduced generalization from large batches. Frequent synchronization of models or gradients among workers leads to significant communication overheads, exacerbated as worker counts and model sizes increase [71, 48, 79]. This challenge intensifies with workers distributed across nodes with limited inter-node communication speeds. To mitigate this, communication-efficient distributed optimization methods [43], such as local gradient methods [67, 82, 81, 26, 73] and their variants, have been developed. In local stochastic gradient methods (Local SGD), workers train locally and synchronize less frequently by averaging models after several local steps, effectively reducing communication overheads. In data-parallel optimization, the total batch size is the sum of all workers’ local batch sizes, increasing with more workers. Large batch sizes are preferred to fully utilize worker memory, making large-batch training the standard in large-scale model training. However, this approach often results in reduced model generalization, a problem noted by LeCun et al. [47]. Large-batch training typically creates a loss landscape with sharp minima, worsening generalization [35]. Although increasing the learning rate proportionally to batch size can mitigate the generalization gap [29, 65, 66], it cannot fully close it. Keskar et al. [35] recommended adaptive sampling to gradually increase batch sizes during training [7, 19], explored in [11, 12, 44, 59]. However, this method is still confined to single- or few-worker scenarios and does not scale well with larger models and more workers.
Motivated by the above two challenges, we answer the following question affirmatively:
Can we develop training strategies which are communication-efficient and memory-efficient with desirable generalization performance for large-scale distributed training?
Contributions.
We make the following contributions: (i) We develop adaptive batch size strategies for local gradient methods by extending the adaptive sampling methods in [7, 3]. Such strategies are for local batch sizes, i.e., batch sizes of individual workers. This is particularly relevant when the workers are heterogeneous devices with different computational speeds and memories, as well as different local objectives and datasets. The proposed strategies are communication-efficient by leveraging local gradient methods, and memory-efficient by leveraging large batch sizes at later stages of training. (ii) Under various standard assumptions, we provide convergence guarantees of the proposed approaches for smooth (strongly) convex and nonconvex objectives under a homogeneous data setting. (iii) We empirically demonstrate the efficacy of the proposed strategies with a three-layer convolutional neural network and ResNet for image classification on the MNIST and CIFAR-10 datasets, respectively, in terms of communication efficiency, memory efficiency, and model generalization performance. To grasp a thorough understanding of the strategies, we also study the effect of different hyperparameters such as the number of local steps and the one which controls the probability of increasing batch sizes.
2 Related Work
Minibatch gradient methods.
Minibatch SGD [13, 63, 4, 15] is the simplest and most widely used optimizer for large-scale machine learning [5], and is easily extendable with data parallelism [49]. The convergence of minibatch SGD and its variants like AdaGrad [18] and Adam [39] is well-understood across various settings, including convex [21], nonconvex [36], and nonsmooth [10] objectives, under different gradient variance assumptions. However, frequent gradient synchronization among workers leads to communication overheads, posing a significant challenge in distributed training.
Local gradient methods.
Local SGD [67, 83, 54], a communication-efficient variant of minibatch SGD, reduces overheads from frequent synchronization in distributed training. Recent studies [51, 17] highlight its efficacy in pre-training large language models with extensive computational resources. Optimization of Local SGD for varied data settings has shown promising convergence rates [67, 77, 76, 40, 37, 69, 74]. It includes extensions like asynchronous SGD [55, 41, 57, 45, 30] and hierarchical SGD [50, 8, 75]. Federated learning [54, 9, 33, 34, 62], another model focusing on efficiency, addresses heterogeneity and partial participation. While applicable, we defer exploring federated learning’s theoretical and empirical aspects to future work.
Generalization gap in large-batch training.
Large-batch training not only faces communication overheads in minibatch SGD under distributed data parallelism but also shows poorer generalization despite better hardware efficiency. It has been successful in tasks like ImageNet classification and BERT training [80, 22, 64]. Addressing these, Lin et al. [50] introduced Post-local SGD, a hybrid of minibatch and Local SGD, where early training involves frequent synchronization of local models, followed by Local SGD. In large-scale environments, Ortiz et al. [58] noted that Post-local SGD does not consistently outperform Local SGD in generalization, and the transition from minibatch to Local SGD presents a trade-off between communication efficiency and generalization. Following these findings, Gu et al. [25, 24] analyzed Local SGD’s generalization via an SDE approximation and proposed the Quadratic Synchronization Rule (QSR) for adjusting local gradient steps during training.
Adaptive batch size strategies.
In stochastic optimization, adaptive sampling methods [7, 19, 3] adjust batch sizes based on gradient noise, enhancing training processes for smooth finite-sum problems. These methods, further explored in deep learning [11, 12, 44, 59], have not been applied to data parallelism with distributed batches. The strategies discussed here draw from batch size scheduling, a key technique in pretraining LLMs, evidenced by its use in GPT-3 [6], Nemotron-4 [60], OLMo 65B [23], and DeepSeek-V2 [14], which involve increasing batch sizes in set stages. Despite their benefits for efficiency and parallelization, these strategies are heuristic and have unclear impacts on training. This work aims to develop theoretically grounded, communication-efficient adaptive batch size strategies akin to adaptive sampling, viewed as stochastic variance-reduction methods [31, 78, 20], focusing on scalable batch size increments.
Several studies have explored adaptive batch size strategies in deep learning, including Big Batch SGD [11, 12], CABS [1], AdaBatch [16], and SimiGrad [61]. Conversely, AdaScale SGD [32] modifies learning rates for large batches instead of batch sizes. These methods, however, are often not theoretically robust, lacking solid convergence proofs or are restricted to SGD analyses under strict conditions like convexity or the Polyak–Łojasiewicz condition. Their batch size adjustment strategies usually depend on heuristic rules, such as geometric growth or decay [61], which can lead to unreliable outcomes. Crucially, these methods overlook the distributed data-parallel context, prevalent in practice, thus not addressing the needs of adaptive batch size strategies in large-scale distributed training.
Hyperparameter tuning for Local SGD.
Besides batch sizes, recent research has also focused on optimizing learning rate schedules for Local SGD, particularly in identical data settings as discussed in [2]. This optimization is crucial in federated learning with heterogeneous data, where local objectives differ significantly in geometry and smoothness, necessitating locally adaptive learning rate schedules [56, 38]. Our adaptive local batch size strategies complement these efforts.
3 Preliminaries
Notation.
[…]
3.1 Problem Formulation
We consider a distributed training setting with $M$ workers, each with a possibly heterogeneous (i.e. non-i.i.d. ) underlying true data distribution $\mathbb{P}_{m}$ ($m\in\llbracket M\rrbracket$). The true data distribution $\mathbb{P}_{m}$ is usually unknown, but can be approximated by its empirical distribution ^$\widehat{\mathbb{P}}_{m}=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}\updelta_{\xi_{i}^{m}}$, where $\mathcal{D}_{m}\coloneqq\{\xi_{i}^{m}\}_{i\in\llbracket n_{m}\rrbracket}$ is the set of $n_{m}$ data samples of worker $m$ and $\xi_{i}^{m}\in\mathcal{Z}_{m}\subseteq\mathbb{R}^{p}$ for $i\in\llbracket n_{m}\rrbracket$. The goal is to find an approximate minimizer of the global objective $F$ which is the average of the local objectives $F_{m}$:
$\operatorname*{minimize}_{x\in\mathbb{R}^{d}}\ F(x)\coloneqq\frac{1}{M}\sum_{m=1}^{M}F_{m}(x),$
where $F_{m}(x)\coloneqq\mathbb{E}_{\xi\sim\widehat{\mathbb{P}}_{m}}[f_{m}(x;\xi)]=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}f_{m}(x;\xi_{i}^{m})$ is the local objective for worker $m$.
At each iteration $k\in\mathbb{N}$, since the number of samples of each worker $n_{m}$ could be large, the per-worker gradient $\nabla F_{m}$ cannot be computed. It can however be approximated by its minibatch counterpart
$(\forall x\in\mathbb{R}^{d})\quad\nabla F_{\mathcal{B}_{k}^{m}}(x)\coloneqq\frac{1}{b_{k}^{m}}\sum_{i\in\mathcal{B}_{k}^{m}}\nabla f_{m}(x;\xi_{i}^{m}),$ (1)
where $\mathcal{B}_{k}^{m}$ is the local batch of worker $m$ at iteration $k$ and $b_{k}^{m}\coloneqq|\mathcal{B}_{k}^{m}|$ is the corresponding local batch size, assuming that $f_{m}(\cdot;\xi^{m})$ is continuously differentiable for any $\xi^{m}\in\mathcal{Z}_{m}$.
3.2 Minibatch and Local Stochastic Gradient Methods
We present both the formulations of minibatch SGD under data parallelism and Local SGD.
Minibatch SGD.
Minibatch SGD under data parallelism is a fully synchronized stochastic gradient method. In particular, at each iteration $k$, after each worker $m$ has computed its local batch gradient $\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$, the global batch gradient $\nabla F_{\mathcal{B}_{k}}(x_{k})$ is computed by averaging all local batch gradients by $\nabla F_{\mathcal{B}_{k}}(x_{k})=\frac{1}{M}\sum_{m=1}^{M}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$ with a global batch size $b_{k}\coloneqq|\mathcal{B}_{k}|=\sum_{m=1}^{M}b_{k}^{m}=Mb_{k}^{1}$, assuming that all local batch sizes are equal. This is followed by the global model update, performed via $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$, where $\alpha_{k}$ is the learning rate of iteration $k$. The same mechanism also applies to other optimizers such as Adam [39] and AdaGrad [18].
Local SGD.
As opposed to minibatch SGD, Local SGD [67] reduces the communication frequency. Communication for model averaging (through all-reduce operations) is performed every $H$ local gradient steps, Local SGD takes the following updates:
$(\forall m\in\llbracket M\rrbracket)(\forall k\in\mathbb{N})\quad x_{k+1}^{m}=\begin{cases}\frac{1}{M}\sum_{m=1}^{M}\left(x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right)&\text{if }H\text{ divides }k+1,\\ x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})&\text{otherwise}.\end{cases}$ (2)
We also consider the following alternative representation that explicitly separates the numbers of local gradient steps $H$ and communication rounds $K$. In particular, for $M$ parallel workers, $K$ rounds of communication, $H$ local gradient steps per round, Local SGD updates can also be expressed as
$(\forall m\in\llbracket M\rrbracket)(\forall k\in\llbracket 0,K-1\rrbracket)\;\begin{cases}x_{k,h+1}^{m}=x_{k,h}^{m}-\alpha_{k,h}\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})&\text{for }h\in\llbracket 0,H-1\rrbracket,\\ x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}.\end{cases}$ (3)
4 Adaptive Local Batch Size Strategies for Local Gradient Methods
The proposed adaptive local batch size strategies are based on the adaptive sampling method [7] originally developed for unconstrained stochastic optimization problems using minibatch SGD. We outline its mechanism and its extension to Local SGD below.
4.1 Adaptive Sampling Methods
Introduced in [7], the adaptive sampling method for minibatch SGD was designed for a single worker ($M=1$) with the objective $F(x)=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i})$ over samples $\mathcal{D}=\{\xi_{i}\}_{i=1}^{n}$. This method relies on a key stochastic optimization principle: when the sample size $n$ is large, the batch gradient $\nabla F_{\mathcal{B}_{k}}$ approximates the full gradient $\nabla F$. If $f(\cdot;\xi)$ is continuously differentiable and convex for any $\xi\in\mathcal{Z}\subseteq\mathbb{R}^{p}$, then $-\nabla F_{\mathcal{B}_{k}}(x_{k})$ serves as a descent direction for $F$ at $x_{k}\in\mathbb{R}^{d}$, provided there exists $\eta\in\left(0,1\right)$ such that
$\delta_{\mathcal{B}_{k}}(x_{k})\coloneqq\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|\leqslant\eta\|\nabla F(x_{k})\|.$ (4)
Note that $\nabla F(x_{k})$ is not available in minibatch SGD, but $\delta_{\mathcal{B}_{k}}(x_{k})$ can be approximated by $\widehat{\delta}_{\mathcal{B}_{k}}(x_{k})^{2}\coloneqq\frac{1}{b_{k}}\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))\cdot\frac{n-b_{k}}{n-1}$, where, for any vector-valued function $h\colon\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}^{d}$, the batch variance is defined by
$\mathrm{Var}_{i\in\mathcal{B}_{k}}(h(x_{k};\xi_{i}))\coloneqq\frac{1}{b_{k}-1}\sum_{i\in\mathcal{B}_{k}}\left\lVert h(x_{k};\xi_{i})-\mathbb{E}_{i\in\mathcal{B}_{k}}[h(x_{k};\xi_{i})]\right\rVert^{2},$ (5)
with $\mathbb{E}_{i\in\mathcal{B}_{k}}[h(x_{k};\xi_{i})]\coloneqq\frac{1}{b_{k}}\sum_{i\in\mathcal{B}_{k}}h(x_{k};\xi_{i})$. Hence, as $n\to\infty$, condition (4) can be approximated as
$\frac{1}{b_{k}}\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))\leqslant\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}.$ (6)
In practice, for each iteration $k\in\mathbb{N}$, the dynamic sample gradient algorithm [7] performs the following minibatch SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with learning rate $\alpha_{k}>0$, and then checks the following (approximate) norm test condition (6), which can also be viewed as an approximation of the exact variance norm test in the stochastic setting:
$\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|^{2}\right]\leqslant\eta^{2}\|\nabla F(x_{k})\|^{2},$ (7)
i.e., condition (4) holds in expectation. Here we abbreviate the conditional expectation on $\mathcal{F}_{k}$ (i.e., the filtration $\mathcal{F}_{k}\coloneqq\sigma(\{x_{0},\mathcal{B}_{0},\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}\})$) by $\mathbb{E}_{k}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$. If condition (6) is violated, the batch size of the next iteration $b_{k+1}$ is increased, determined by the formula:
$b_{k+1}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right\rceil,$
and the next batch $\mathcal{B}_{k+1}$ is sampled with this larger batch size accordingly. Otherwise, the next batch size remains the same as the current one. The value of the pre-specified constant $\eta$ controls the probability of obtaining a descent direction, thereby affecting the probability of increasing the batch size as well as the magnitude of the next batch size.
Condition (7) is more relaxed than the typical assumption of uniformly bounded gradient variance for minibatch stochastic gradient methods: for any i.i.d.  batch $\mathcal{B}\subset\llbracket n\rrbracket$ and some positive constant $\sigma>0$, $\mathbb{E}_{\mathcal{B}}\left[\|\nabla F_{\mathcal{B}}(x)-\nabla F(x)\|^{2}\right]\leqslant\sigma^{2}$ for all $x\in\mathbb{R}^{d}$, which is usually unverifiable in practice. The adaptive batch size schedule using the approximate norm test may cause rapid batch size increases, undermining its purpose. To moderate these schedules, Bollapragada et al. [3] introduced the (augmented) inner product test, which manages the variance between the batch and full gradient’s inner product. Given the preference for large batch sizes for memory efficiency, controlled via $\ eta$, we focus on the norm test and defer exploring local variants of the (augmented) inner product test to future work.
4.2 Extensions to Local Gradient Methods
We extend the above adaptive sampling methods beyond the single worker setting with data parallelism using multiple workers, more specifically for Local SGD. We have to invoke a local variant of the norm test (7), called the (exact variance) local norm test, at every iteration $k\in\mathbb{N}$, using formulation (2): for some sequence of positive constants $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$,
$(\forall m\in\llbracket M\rrbracket)\quad\mathbb{E}_{\mathcal{F}_{k}^{m}}\left[\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})-\nabla F_{m}(x_{k}^{m})\|^{2}\right]\leqslant\eta_{m}^{2}\left\lVert\nabla F_{m}(x_{k}^{m})\right\rVert^{2},$ (8)
where $\mathbb{E}_{\mathcal{F}_{k}^{m}}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}^{m}]$ denotes the conditional expectation on $\mathcal{F}_{k}^{m}\coloneqq\sigma(\{x_{0},\mathcal{B}_{0}^{m},\ldots,\mathcal{B}_{k-1}^{m}\})$. Again, $\nabla F_{m}$ is in general unavailable since the sample size $n_{m}$ could be large. The above local norm test is approximately enforced by the following (approximate) local norm test:
$(\forall m\in\llbracket M\rrbracket)\quad\frac{1}{b_{k}^{m}}\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}\left(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m})\right)\leqslant\eta_{m}^{2}\left\lVert\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right\rVert^{2}.$ (9)
Should the condition (9) not be met, the subsequent local batch sizes are determined by the formula:
$(\forall m\in\llbracket M\rrbracket)\quad b_{k+1}^{m}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m}))}{\eta_{m}^{2}\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\|^{2}}\right\rceil,$ (10)
and the local batches $\mathcal{B}_{k+1}^{m}$ are augmented accordingly. In a practical implementation, for simplicity, we can set $\eta_{m}\equiv\eta$ for all $m\in\llbracket M\rrbracket$ and choose the next local batch sizes to be
$(\forall m\in\llbracket M\rrbracket)\quad b_{k+1}^{m}\equiv b_{k+1}\coloneqq\max_{m\in\llbracket M\rrbracket}\ \left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\|^{2}}\right\rceil,$ (11)
which is particularly useful when homogeneous workers and identical data on the workers are used, so that there would not be discrepancy of training time for the local steps due to different local batch sizes and undesirable stragglers before model averaging.
The proposed adaptive batch size strategies extend to local variants of minibatch stochastic gradient optimizers beyond SGD [63], including momentum SGD [70], AdaGrad [18], Adam [39], and AdamW [52]. The algorithm is detailed in Algorithm 1.
In a data parallel setting, adaptive global batch size strategies can be developed using minibatch SGD, leading to increased communication overheads from aggregating local batch gradient variances among workers to compute the quantities in (6).
Algorithm 1 Adaptive Local Batch Size Schedules for Local Gradient Methods
no. of workers $M\in\mathbb{N}^{*}$, no. of communication steps $K\in\mathbb{N}^{*}$, no. of local steps $H\in\mathbb{N}^{*}$, no. of training samples $N\in\mathbb{N}^{*}$, global step counter $k=0$, processed samples counter $B=0$, training sets $(\mathcal{D}_{m})_{m\in\llbracket M\rrbracket}=(\{\xi_{i}^{m}\}_{i=1}^{n_{m}})_{m\in\llbracket M\rrbracket}$, initial model parameters $x_{0,0}^{m}=x_{0}\in\mathbb{R}^{d}$for all $m\in\llbracket M\rrbracket$, initial local batch sizes $(b_{0}^{m})_{m\in\llbracket M\rrbracket}$, learning rate schedules $(\alpha_{k,h})_{k\in\llbracket 0,K-1\rrbracket,h\in\llbracket 0,H-1\rrbracket}$, $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$$B<N$worker $m=1,\ldots,M$in parallel local step $h=0,\ldots,H-1$
Sample the local batch $\mathcal{B}_{k,h}^{m}$ of size $b_{k,h}^{m}\coloneqq|\mathcal{B}_{k,h}^{m}|$ uniformly from $\mathcal{D}_{m}$
Compute per-sample gradients $\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$ for each $i\in\mathcal{B}_{k,h}^{m}$
Compute the local batch gradient $\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})=\frac{1}{b_{k,h}^{m}}\sum_{i\in\mathcal{B}_{k,h}^{m}}\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$
Compute $\mathrm{Var}_{i\in\mathcal{B}_{k,h}^{m}}(\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m}))$ by applying (5)
Compute the local norm test statistic $\mathsf{T}_{k,h}^{m}$ defined by (10)
Determine the next local batch sizes: $b_{k,h+1}^{m}=\max\{\mathsf{T}_{k,h}^{m},b_{k,h}^{m}\}$
[…]
Update the global model: $x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}$ for each $m\in\llbracket M\rrbracket$triangleright all-reduce
$B\leftarrow B+\sum_{m=1}^{M}\sum_{h=0}^{H-1}b_{k,h}^{m}$
[…]
5 Convergence Analysis
Our analysis focuses on the i.i.d.  setting as shown in (1). We set $\eta_{m}\equiv\eta\in(0,1)$ for all $m\in\llbracket M\rrbracket$, typical in datacenter environments for distributed training [13], where workers access the entire data set. The analysis of heterogeneous data settings is deferred to future work. Our convergence analysis, following [69], does not assume bounded gradient variance across $\mathbb{R}^{d}$. Instead, our adaptive local batch size strategies ensure the expected strong growth condition [36, 5] at each local iterate $x_{k}^{m}$ for all workers $m\in\llbracket M\rrbracket$. This approach relaxes the standard assumptions of uniform boundedness and expected strong growth conditions [72], which are challenging to verify except in the interpolation regime [53].
We might need to invoke the following assumptions on the objective function $F$.
Assumption 1 ($L$-Lipschitz smoothness).
The function $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable, bounded below by $F^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}F(x)=F(x^{\star})\in\mathbb{R}$ with $x^{\star}=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}F(x)$, and is $L$-Lipschitz smooth for some $L>0$, i.e., for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $\left\lVert\nabla F(x)-\nabla F(y)\right\rVert\leqslant L\left\lVert x-y\right\rVert$.
Assumption 2 ($\mu$-strong convexity).
The function $F$ is $\mu$-strongly convex for some $\mu\geqslant 0$, i.e., for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $F(x)-F(y)+\frac{\mu}{2}\left\lVert x-y\right\rVert^{2}\leqslant\langle\nabla F(x),x-y\rangle$. If $\mu=0$, we say that $F$ is convex.
We derive the following convergence results for the proposed adaptive batch size strategies as follows. Their proofs are deferred to Appendix A.
Theorem 1 (Strongly convex; $\mu>0$).
Suppose that Assumptions 1 and 2 hold with $\mu>0$. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$ and a constant learning rate $\alpha_{k}\equiv\alpha$. Then, for some $\alpha\leqslant 1/(10L(HM+\eta^{2}))$, we have
$\mathbb{E}[F(x_{\mathsf{out}})]-F^{\star}=\widetilde{\mathscr{O}}\left(L(HM+\eta^{2})\exp\left\{-\frac{\mu K}{10L(HM+\eta^{2})}\right\}\left\lVert x_{0}-x^{\star}\right\rVert^{2}\right),$ (12)
where $x_{\mathsf{out}}\in\{x_{k}^{m}\}_{k\in\llbracket 0,K-1\rrbracket,m\in\llbracket M\rrbracket}$ is chosen as $x_{k}^{m}$ with probability proportional to $(1-\mu\alpha/2)^{-k}$ uniformly on $m\in\llbracket M\rrbracket$ and $\widetilde{\mathscr{O}}$ hides logarithmic factors.
Theorem 2 (Convex; $\mu=0$).
Suppose that Assumptions 1 and 2 hold with $\mu=0$. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$ and a constant learning rate $\alpha_{k}\equiv\alpha$. Then, for some $\alpha\leqslant 1/(10L(HM+\eta^{2}))$, we have
$\mathbb{E}[F(x_{\mathsf{out}})]-F^{\star}=\mathscr{O}\left(\frac{L(HM+\eta^{2})}{K}\left\lVert x_{0}-x^{\star}\right\rVert^{2}\right),$ (13)
where $x_{\mathsf{out}}$ is chosen uniformly at random from $\{x_{k}^{m}\}_{k\in\llbracket 0,K-1\rrbracket,m\in\llbracket M\rrbracket}$.
Theorem 3 (Nonconvex).
Suppose that Assumption 1 holds. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$ and a constant learning rate $\alpha_{k}\equiv\alpha$. Then, for some $\alpha\leqslant 1/(10L(HM+\eta^{2}))$, we have
$\mathbb{E}\left\lVert\nabla F(x_{\mathsf{out}})\right\rVert^{2}=\mathscr{O}\left(\frac{L(HM+\eta^{2})}{K}(F(x_{0})-F^{\star})\right),$ (14)
where $x_{\mathsf{out}}$ is chosen uniformly at random from $\{x_{k}^{m}\}_{k\in\llbracket 0,K-1\rrbracket,m\in\llbracket M\rrbracket}$.
The theorems confirm that Local SGD, using the local norm test with exact variance, achieves linear convergence rates for both convex and nonconvex objectives. This scaling is proportional to the number of local gradient steps $H$ and workers $M$, akin to Local SGD under constant batch sizes and bounded gradient variance [69].
Remark 1 (Choice of $\eta$).
Convergence rates depend quadratically on $\eta$, which controls the likelihood and size of increased local batch sizes. Choosing smaller $\eta$ values accelerates convergence but causes quick batch size escalation, undermining adaptive batch sizing and widening the generalization gap.
6 Numerical Experiments
We evaluate the efficiency of the training (total steps or wall-clock time) and the generalization performance of our strategies through image classification tasks. We use a three-layer convolutional neural network on the MNIST dataset [46] and a ResNet-50 [28] on the CIFAR-10 dataset [42], with four workers ($M=4$ GPUs). The models are trained using Local Adam and Local SGD with momentum (SHB), synchronizing only the models every $H$ steps. Further details and results are in Appendix B.
6.1 Three-layer CNN on MNIST
We demonstrate our strategies using a three-layer CNN for MNIST classification. The training involves 12 million samples (200 epochs) with global and local batch sizes set to 60,000 and 15,000, respectively. This task uses a fixed learning rate of 0.0008 without any rate schedules. Results are in Table 1.
Table 1: Adaptive local batch size strategies for three-layer CNN on MNIST; no. of total steps (steps), wall-clock time (in minutes), average local batch sizes (bsz.), and top validation accuracy (acc.; in %)
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 1456 5.01 2048 99.10 1464 9.44 2048 99.06 1464 25.58 2048 99.08
Constant 352 6.84 8192 98.80 364 8.19 8192 98.85 366 12.37 8192 98.94
Constant 192 10.59 15000 98.40 196 11.41 15000 98.43 199 13.88 15000 98.54
$\eta=0.25$ 213 11.35 13414 98.51 221 11.94 13427 98.55 223 14.50 13467 98.66
$\eta=0.375$ 224 11.85 13077 98.61 228 12.16 13031 98.58 228 14.86 13078 98.71
$\eta=0.50$ 229 11.85 12615 98.58 235 11.73 12648 98.60 235 14.42 12748 98.72
Figure 1: Validation accuracy and local batch sizes of Local Adam with adaptive batch size strategies ($H=16$).
Values of $\eta$.
Table 1 shows that adaptive local batch size strategies can match or exceed the validation accuracy of constant batch size strategies. Specifically, with 16 local gradient steps and $\eta=0.25$, the adaptive strategy achieves an average batch size similar to a constant batch size of 15,000 (see Figure 1), but with superior accuracy and 21 additional steps. This confirms that adaptive strategies can mitigate the generalization gap in large-batch training. Furthermore, Figure 2 indicates that these strategies reach a maximum local batch size of 15,000 later in the training, enhancing hardware utilization and efficiency. Similar trends are noted for $H=4$ and $H=1$.
Figure 2: Validation accuracy and local batch sizes of Local Adam with adaptive batch size strategies with $\eta\in\{0.25,0.375,0.5\}$ and $H\in\{16,4,1\}$.
Local gradient steps $H$.
Smaller $H$ increases training time due to more frequent communication, and contrary to expectations, does not always enhance generalization. As Table 1 demonstrates, with a local batch size of 2048, $H=1$ and $H=4$ have lower accuracy compared to $H=16$. This pattern is consistent with adaptive batch sizes, where $H=4$ underperforms $H=16$ at $\eta=0.375$. Moreover, while $H=1$ leads to longer training times, it achieves higher accuracy.
6.2 ResNet-50 on CIFAR-10
We scale our experiment to ResNet-50 for CIFAR-10, using a training budget of 12 million samples (200 epochs) and setting maximum global and local batch sizes of 60,000 and 15,000, respectively. Our learning rate schedules include a 10% linear warmup and cosine decay, with peak and minimum rates of 0.5 and 0.05, respectively.
Table 2: Adaptive local batch size strategies for ResNet-50 on CIFAR-10; no. of total steps (steps), wall-clock time (in minutes), average local batch sizes (bsz.), and top validation accuracy (acc.; in %)
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 2432 27.16 1024 92.31 2440 34.00 1024 92.65 2441 60.51 1024 91.11
Constant 608 34.99 4096 63.50 608 37.10 4096 57.04 610 46.39 4096 40.40
Constant 304 34.31 8192 12.51 304 35.34 8192 16.61 305 40.01 8192 14.78
$\eta=0.8$ 576 90.73 4306 81.95 416 138.39 6001 75.09 241 50.67 10380 25.32
$\eta=0.85$ 512 92.41 4900 81.95 912 141.51 2742 86.89 433 144.75 5767 72.66
$\eta=0.9$ 2816 96.98 885 93.59 1982 126.38 1261 92.53 3940 111.92 634 93.75
Values of $\eta$.
Table 3 shows that adaptive local batch size strategies can achieve comparable or superior validation accuracy to constant batch sizes with similar averages. For instance, with $H=16$ and $\eta=0.80$, the adaptive strategy reaches an average batch size akin to a constant batch size of 4096 but with nearly 50% higher accuracy and 32 fewer steps, illustrating the effectiveness of adaptive strategies in bridging the generalization gap in large-batch training.
Figure 3: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies ($H=16$).
Local gradient steps $H$.
We first examine constant batch sizes (Table 3 and Figure 4). Local gradient methods, as noted in [50, 58, 24, 25], may improve generalization compared to synchronized methods ($H=1$), although we employ SHB in synchronized models instead of standard SGD. For adaptive local batch sizes, smaller $H$ values increase batch sizes more rapidly (Table 3), supporting our theory that a smaller $\eta$ should be used with a larger $H$ to maintain convergence rates of $\mathscr{O}(HM+\eta^{2})$. Refer to Figures 3 and B.2 for further illustrations.
Figure 4: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies with $\eta\in\{0.25,0.375,0.5\}$ and $H\in\{16,4,1\}$.
Number of workers $M$.
We conducted additional experiments with $M=8$. Compared to the 4-worker scenario, increasing the number of workers reduces training time but lowers average accuracy, indicating poorer generalization. Nonetheless, the observations regarding local gradient steps $H$ and values of $\eta$ remain consistent with the 4-worker case.
Figure 5: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies ($H=16$) with 8 workers.
Table 3: Adaptive local batch size strategies for ResNet-50 on CIFAR-10 with 8 workers
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 2432 12.06 512 92.06 2440 15.45 512 92.87 2441 29.09 512 92.41
Constant 1216 11.19 1024 88.88 1220 12.93 1024 87.87 1220 19.72 1024 86.80
Constant 608 10.72 2048 66.94 608 11.59 2048 71.32 610 15.03 2048 59.57
$\eta=0.8$ 213 11.55 5666 17.88 216 11.82 5655 15.98 218 12.29 5650 16.84
$\eta=0.85$ 220 12.22 5610 19.47 222 12.30 5570 22.06 222 12.66 5569 21.99
$\eta=0.9$ 261 15.25 5236 28.78 277 16.92 5052 33.07 278 18.26 4959 36.83
7 Concluding Remarks
This work introduces adaptive local batch size strategies for local gradient methods in data parallelism, focusing on reducing communication overheads by increasing local batch sizes independently. Although our analysis and experiments are limited to identical data settings, the algorithm is applicable to heterogeneous scenarios. Future work will explore convergence in diverse settings and federated learning [54, 33], including personalized approaches [27]. Extensions to asynchronous SGD [55, 41, 57, 45, 30] and hierarchical SGD [50, 8, 75] are also promising for improving communication efficiency and addressing worker heterogeneity.
Acknowledgments
The research of Han Liu is supported by NIH R01LM01372201, NSF RI 1840857, NSF TRIPOD 1740735, NSF DMS1454377-CAREER, NSF IIS 1546482, along with an Alfred P. Sloan Fellowship. The research of Mladen Kolar is supported in part by NSF ECCS-2216912. This research is supported in part through the computational resources and staff contributions from the Data Science Institute at the University of Chicago, through its AI + Science Research Funding Initiatives.
Appendix
Appendix A Proofs of Main Text
We present the omitted proofs of the main text, mainly by simplifying the technique in [69].
Notation.
Let us define
$(\forall k\in\mathbb{N})\quad\overline{x}_{k+1}\coloneqq\overline{x}_{k}-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}g_{k}^{m},$ (A.1)
where $g_{k}^{m}\coloneqq\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})$ and $\overline{x}_{0}\coloneqq x_{0}$.
Let use recall that $\mathbb{E}_{k}\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$, $\mathbb{E}_{\mathcal{F}_{k}^{m}}\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}^{m}]$, and $\mathbb{E}$ denotes the unconditional expectation.
We also define the notion of slowly decreasing learning rates, called $\tau$-slow sequences.
Definition A.1 ($\tau$-slow sequence; Definition 10 in in [69]).
A sequence $(a_{k})_{k\in\mathbb{N}}\subset(0,\infty)$ with positive values is $\tau$-slow decreasing for $\tau\geqslant 1$ if
$(\forall k\in\mathbb{N})\quad a_{k+1}\leqslant a_{k}\qquad\text{and}\qquad a_{k+1}\left(1+\frac{1}{2\tau}\right)\geqslant a_{k}.$
The sequence $(a_{k})_{k\in\mathbb{N}}$ is $\tau$-slow increasing if $(1/a_{k})_{k\in\mathbb{N}}$ is $\tau$-slow decreasing.
A.1 Technical Lemmas
We first state without proofs some preparatory lemmas, which are either standard results or whose proofs are be found in the cited work.
Lemma A.1.
For any $M$ vectors $v_{1},\ldots,v_{M}\in\mathbb{R}^{d}$, we have
$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert\leqslant\sum_{m=1}^{M}\left\lVert v_{m}\right\rVert.$ (A.2)
Lemma A.2.
For any $M$ vectors $v_{1},\ldots,v_{M}\in\mathbb{R}^{d}$, we have
$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert^{2}\leqslant M\sum_{m=1}^{M}\left\lVert w_{m}\right\rVert^{2}.$ (A.3)
Lemma A.3.
For any two vectors $x,y\in\mathbb{R}^{d}$ and any positive number $\rho>0$, we have
$\left\lVert x+y\right\rVert^{2}\leqslant(1+\rho)\left\lVert a\right\rVert^{2}+(1+1/\rho)\left\lVert b\right\rVert^{2}.$ (A.4)
Lemma A.4 (Cauchy–Schwarz inequality).
For any two vectors $x,y\in\mathbb{R}^{d}$ and any positive number $\rho>0$, we have
$2\langle x,y\rangle\leqslant\rho\left\lVert a\right\rVert^{2}+\rho^{-1}\left\lVert b\right\rVert^{2}.$ (A.5)
Lemma A.5.
If $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is $L$-Lipschitz smooth for some $L>0$ (Assumption 1), then we have
$(\forall x\in\mathbb{R}^{d})\quad\left\lVert\nabla F(x)\right\rVert^{2}\leqslant 2L(F(x)-F^{\star}).$ (A.6)
Lemma A.6 (Lemma 2 in [68]; Lemma 13 in [69]).
For any nonnegative sequence $(u_{k})_{k\in\mathbb{N}}$ and parameters $\kappa\geqslant\chi>0$, $\sigma\geqslant 0$ and $K\in\mathbb{N}$, there exists a constant $\alpha\leqslant 1/\kappa$ such that for constant learning rates $(\alpha_{k}\equiv\alpha)_{k\in\mathbb{N}}$ and weights $w_{k}\coloneqq(1-\chi\alpha)^{-k-1}$, we have
$\Psi_{K}\coloneqq\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{(1-\chi\alpha_{k})u_{k}}{\alpha_{k}}-\frac{u_{k+1}}{\alpha_{k}}+\sigma\alpha_{k}\right)=\widetilde{\mathscr{O}}\left(\kappa u_{0}\mathrm{e}^{-\chi K/\kappa}+\frac{\sigma}{\chi K}\right),$
where $W_{K}\coloneqq\sum_{k=1}^{K}w_{k}$.
Lemma A.7 (Lemma 14 in [69]).
For any nonnegative sequence $(u_{k})_{k\in\mathbb{N}}$ and parameters $\kappa\geqslant 0$, $\sigma\geqslant 0$ and $K\in\mathbb{N}$, there exists a constant $\alpha\leqslant 1/\kappa$ such that for constant learning rates $(\alpha_{k}\equiv\alpha)_{k\in\mathbb{N}}$, we have
$\Psi_{K}\coloneqq\frac{1}{K+1}\sum_{k=0}^{K}\left(\frac{u_{k}}{\alpha_{k}}-\frac{u_{k+1}}{\alpha_{k}}+\sigma\alpha_{k}\right)\leqslant\frac{\kappa u_{0}}{K+1}+2\sqrt{\frac{\sigma u_{0}}{K+1}}.$
A.2 Proof of Theorems 1, 2 and 3
We first prove a descent lemma for ($\mu$-strongly) convex objectives.
Lemma A.8 (Descent lemma for ($\mu$-strongly) convex objectives).
Suppose that Assumptions 1 and 2 hold with $\mu>0$. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$. If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(4L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence$(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (A.1) satisfies
$\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\leqslant\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{\alpha_{k}}{2M}\sum_{m=1}^{M}\mathbb{E}[F(x_{k}^{m})-F^{\star}]+\frac{3L\alpha_{k}}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
Proof.
The proof is similar to that of Lemma 25 in [69]. We state it here for convenience and completeness. Expanding we have
$\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}=\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle g_{k}^{m},x_{k}^{m}-x^{\star}\rangle+\frac{\alpha_{k}^{2}}{M^{2}}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}+\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle g_{k}^{m},x_{k}^{m}-\overline{x}_{k}\rangle.$
Taking expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\\ \leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle+\frac{\alpha_{k}^{2}}{M^{2}}\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}+\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-\overline{x}_{k}\rangle.$
Note that $\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}$ can be bounded by
$\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}$ $=\left\lVert\sum_{m=1}^{M}\nabla F(x_{k}^{m})\right\rVert^{2}+\sum_{m=1}^{M}\mathbb{E}\left\lVert g_{k}^{m}-\nabla F(x_{k}^{m})\right\rVert^{2}$
$\leqslant M\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\eta^{2}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (A.3) and (8) (A.7)
[…]
On the other hand, by Assumption 2, we also have
$-2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle\leqslant-\mu\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}-2\left(F(x_{k}^{m})-F^{\star}\right).$ (A.8)
By Cauchy–Schwarz inequality (A.5) with $\rho=2L$, we have
$2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-\overline{x}_{k}\rangle$ $\leqslant 2L\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+\frac{1}{2L}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$
[…]
By (A.4), we also have
$-\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}\leqslant-\frac{1}{2}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$ (A.10)
Thus, plugging in (A.6), (A.9) and (A.10), we obtain
$\mathbb{E}_{k}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}$ $\leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle+\frac{2L\alpha_{k}^{2}(M+\eta^{2})}{M^{2}}\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)$
$\qquad+\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-\overline{x}_{k}\rangle$
$\leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\left[\mu\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+2\left(F(x_{k}^{m})-F^{\star}\right)\right]$
$\qquad+\frac{2L(M+\eta^{2})\alpha_{k}^{2}}{M^{2}}\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)+\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\left[2L\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+F(x_{k}^{m})-F^{\star}\right]$
$\leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\frac{\mu\alpha_{k}}{M}\sum_{m=1}^{M}\left[-\frac{1}{2}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\right]$
$\qquad-\frac{\alpha_{k}}{M}\left(1-\frac{2L(M+\eta^{2})\alpha_{k}}{M}\right)\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)+\frac{2L\alpha_{k}}{M}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
$\leqslant\left(1-\frac{\mu\alpha_{k}}{2}\right)\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{\alpha_{k}}{2M}\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)+\frac{3L\alpha_{k}}{M}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2},$
since $\alpha_{k}\leqslant M/(4L(M+\eta^{2}))$ and $\mu\leqslant L$ if $F$ is both $\mu$-strongly convex and $L$-Lipschitz smooth. Taking total expectation gives the desired result. ∎
We also prove a descent lemma for general nonconvex objectives.
Lemma A.9 (Descent lemma for nonconvex objectives).
Suppose that Assumption 1 holds. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$. If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(2L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence$(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (A.1) satisfies
$\mathbb{E}[F(\overline{x}_{k+1})]\leqslant\mathbb{E}[F(\overline{x}_{k})]-\frac{\alpha_{k}}{4M}\sum_{m=1}^{M}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{L^{2}\alpha_{k}}{2M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
Proof.
The proof is similar to that of Lemma 26 in [69]. We state it here for convenience and completeness.
By Assumption 1, we have
$F(\overline{x}_{k+1})\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),g_{k}^{m}\rangle+\frac{L\alpha_{k}^{2}}{2M^{2}}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}.$
Taking expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}[F(\overline{x}_{k+1})]$ $\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}}{2M^{2}}\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}$
$\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (A.7)
$=F(\overline{x}_{k})-\left(\frac{\alpha_{k}}{M}-\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\right)\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$
$\qquad+\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle.$
By Cauchy–Schwarz inequality (A.5) with $\rho=1$ and Assumption 1, we have
$\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle$ $\leqslant\frac{1}{2}\sum_{m=1}^{M}\left[\left\lVert\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k})\right\rVert{\nabla F(x_{k}^{m})}^{2}+\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}\right]$
$\leqslant\frac{L^{2}}{2}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}+\frac{1}{2}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}.$
Plugging this in, we obtain
$\mathbb{E}_{k}[F(\overline{x}_{k+1})]$ $=F(\overline{x}_{k})-\frac{\alpha_{k}}{2M}\left(1-\frac{L(M+\eta^{2})\alpha_{k}}{M}\right)\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{L^{2}\alpha_{k}}{2M}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
$\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{4M}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{L^{2}\alpha_{k}}{2M}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2},$
since $\alpha_{k}\leqslant M/(2L(M+\eta^{2}))$. Taking total expectation gives the desired result. ∎
We then derive an upper bound for $\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$.
Lemma A.10.
Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$, and $(\overline{x}_{k})_{k\in\mathbb{N}}$ be the sequence defined in (A.1). If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant 1/(10L(HM+\eta^{2}))$ for all $k\in\mathbb{N}$ and $(\alpha_{k}^{2})_{k\in\mathbb{N}}$ is $H$-slow decreasing, then we have
$\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{1}{10LHM}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.11)
Moreover, for any $H$-slow increasing nonnegative sequence $(w_{k})_{k\in\mathbb{N}}$ we have
$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$
Proof.
We first consider
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$ $=\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{\lfloor K/H\rfloor H}-(\overline{x}_{k}-\overline{x}_{\lfloor K/H\rfloor H})\right\rVert^{2}$
$\leqslant\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{\lfloor K/H\rfloor H}\right\rVert^{2}$
$\leqslant\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert\sum_{h=0}^{H-1}\alpha_{k-h}\left(\nabla F(x_{k-h}^{m})+g_{k-h}^{m}-\nabla F(x_{k-h}^{m})\right)\right\rVert^{2}$
$\leqslant\frac{3H}{2M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}$
$\qquad\qquad+\frac{3}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert g_{k-h}^{m}-\nabla F(x_{k-h}^{m})\right\rVert^{2}$ by (A.4) and (A.3)
$\leqslant\frac{3(H+\eta^{2})}{2K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$
The first inequality follows from $\mathbb{E}\left\lVert X-\mathbb{E}X\right\rVert^{2}\leqslant\mathbb{E}\left\lVert X\right\rVert^{2}$ for any random variable $X$ in $\mathbb{R}^{d}$. Since $(\alpha_{k}^{2})_{k\in\mathbb{N}}$ is $H$-slow decreasing, for $h\leqslant H$, we have $\alpha_{k-h}^{2}\leqslant\alpha_{k}^{2}(1+1/(2H))^{H}\leqslant\alpha_{k}^{2}\mathrm{e}^{H/(2H)}\leqslant 2\alpha_{k}^{2}$ since $1+x\leqslant\mathrm{e}^{x}$ for any $x\in\mathbb{R}$. Then, we can simply the inequality:
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{3(H+\eta^{2})\alpha_{k}^{2}}{K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.12)
Since $\alpha_{k}\leqslant 1/(10L(HM+\eta^{2}))$ for all $k\in\mathbb{N}$, we have
$3L\cdot 3(H+\eta^{2})\alpha_{k}^{2}\leqslant\frac{1}{10L(HM+\eta^{2})}\leqslant\frac{1}{10LHM}.$
Since $(w_{k})_{k\in\mathbb{N}}$ is $H$-slow increasing, we have $w_{k}\leqslant w_{k-h}(1+2/(2H))^{h}\leqslant w_{k-h}(1+2/(2H))^{H}\leqslant w_{k-h}\sqrt{\mathrm{e}}\leqslant 2w_{k-h}$ for each $h\in\llbracket 0,H\rrbracket$. Therefore, we have
$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$ $\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}\frac{w_{k}}{2H}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}$
$\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}\frac{1}{H}\sum_{h=0}^{H-1}w_{k-h}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}$
$\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.13)
Proof of Theorem 1.
Using Lemma A.10 with $\mu>0$ and (A.6), we have
$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{2}{5M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k-h}^{m})-F^{\star}].$ (A.14)
Since $\alpha_{k}\equiv\alpha\leqslant 1/(10L(HM+\eta^{2}))\leqslant M/(4L(M+\eta^{2}))$, by Lemma A.8, we have
$\frac{1}{2M}\sum_{m=1}^{M}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{1}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{1}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}+\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
Multiplying the weight $w_{k}$ and summing over $k\in\llbracket 0,K\rrbracket$, and by (A.14), we have
$\quad\,\frac{1}{2M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]$
$\leqslant\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right)+\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
$\leqslant\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right)+\frac{2}{5M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k-h}^{m})-F^{\star}],$
which is equivalent to
$\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{10}{W_{K}}\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right).$ (A.15)
We observe that $1-\mu\alpha/2\geqslant 1-\mu/(20L(HM+\eta^{2}))\geqslant 1-1/(8HM)$, we choose $w_{k}=(1-\mu\alpha/2)^{-k-1}$ which are $2H$-slow increasing. Then by Lemma A.6, we obtain the desired result. ∎
Proof of Theorem 2.
For the convex case with $\mu=0$, we simply use Lemma A.6 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$. ∎
Proof of Theorem 3.
By Lemma A.9 and (A.6), and define $u_{k}\coloneqq 4\cdot\mathbb{E}[F(\overline{x}_{k})-F^{\star}]$, we have
$\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ $\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right)+\frac{2L\alpha_{k}}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
$\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right)+\frac{2\alpha_{k}}{15LMW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2},$
by (A.13). We simplify this to
$\left(1-\frac{2\alpha_{k}}{15L}\right)\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right).$
Applying Lemma A.7 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$ gives the desired result.
Appendix B Details of Numerical Experiments
All experiments are conducted on a Slurm cluster. For experiments with 4 GPUs, each node has either 4 NVIDIA GEFORCE RTX 2080 Ti (11GB) GPUs or 4 NVIDIA A40 (48GB) GPUs or 4 NVIDIA L40s (48GB) GPUs. For experiments with 8 GPUs, each node has 8 NVIDIA A100 (80GB) GPUs. For CPUs, each node has either 64 cores or 96 cores. The total RAM is 512GB or 1TB.
B.1 Three-layer Convolutional Neural Network on MNIST
Table 4: Training hyperparameters for three-layer CNN on MNIST
Model Three-layer CNN on MNIST
Training samples 6M (100 epochs)
Weight initialization Default
Optimizer Adam
Learning rate schedule None
$(\beta_{1},\beta_{2})$ $(0.9,0.95)$
$\varepsilon$ $10^{-8}$
Base learning rate 0.0008
Base batch size 8
Maximum batch size 50000
Weight decay 0
Base global batch size 4096
Base local batch size 1024
Maximum global batch size 60000
Maximum local batch size 15000
Weight decay 0
Figure 6: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies with three different seeds.
Figure 7: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=16$) with two other different seeds.
Figure 8: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with three different seeds.
Figure 9: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with three different seeds.
B.2 ResNet-50 on CIFAR-10
Table 5: Training hyperparameters for ResNet-50 on CIFAR-10
Model ResNet-50 on CIFAR-10
Training samples 10M (200 epochs)
Weight initialization Default
Optimizer Momentum SGD (SHB)
Learning rate schedule Linear warmup + cosine decay
Learning rate warmup (samples) 1M (10%)
Peak learning rate 0.05
Minimum learning rate 0.005
Base learning rate 0.05
Base global batch size 256
Base local batch size 64
Maximum global batch size 50000
Maximum local batch size 12500
Weight decay 0.0005
Momentum 0.9
Figure 10: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies with three different seeds.
Figure 11: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies (local gradient steps $H=16$) with two other different seeds.
Figure 12: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with three different seeds.
Figure 13: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with three different seeds.
Figure 14: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies with 8 workers with three different seeds.
Figure 15: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies (local gradient steps $H=16$) with 8 workers with two other different seeds.
Figure 16: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with 8 workers with three different seeds.
Figure 17: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with 8 workers with three different seeds.